OpenStudy (anonymous):

The probability that a typewritten page is free from error is 1/2; the probability that it contains exactly n errors is inversely proportional to n for n =1,2,3,4; and the probability that it contains more than 4 errors is 1/12. Calculate the probability that i) a page will contain exactly one error [Ans: 1/5] ii) six pages will consist of four pages containing no error and the remaining two pages containing exactly one error each [Ans:3/80] I think I must involve the binomial distribution here but i dont know to apply the given information into in to get the answers. Please help.

OpenStudy (anonymous):

As much as i understand: 1/2+ k*1/n +1/12 =1 or by simpifying a bit: k*1/n=5/12 this represents the probability of a page to contain 1,2,3 or 4 errors so for n=1, k=5/12 and the probability is 5/12